OpenStudy (anonymous):

find the value of the function sin 9pi/7

OpenStudy (anonymous):

\[\sin(\frac{ 9 }{ 7 }\pi)= -\sin(\frac{ 2 }{ 7 }\pi)\]

OpenStudy (anonymous):

How are you supposed to find it? Can't you use a calculator? Or a table?

OpenStudy (anonymous):

i need it not in fractions, but decimals :/

OpenStudy (anonymous):

lets assume i know NOTHING about trig and im being asked this question.

OpenStudy (anonymous):

well then we need to start explaining about the unit circle and what sin is

OpenStudy (anonymous):

i know the unit circle i just... dont know how to apply it

OpenStudy (anonymous):

i do it on my calculator and get "-0.7818"

OpenStudy (anonymous):

you need degrees?

OpenStudy (anonymous):

or radians?

OpenStudy (anonymous):

\(\pi\) radians = 180 degrees

OpenStudy (anonymous):

so \[\sin(\frac{ 9 }{ 7 }\pi) \rightarrow \sin(\frac{ 9 }{ 7 } \times 180)\]

OpenStudy (anonymous):

−0.781831482

OpenStudy (anonymous):

but often calculators have an option to use radians instead of degrees

OpenStudy (anonymous):

Your answer was correct